1)     Suppose that the market demand for Blue band is a straight line of the form Q = 300-50p where Q is the quantity bought in thousands of boxes per week and P is the price per box in. 


1
Expert's answer
2021-10-25T09:35:49-0400

"When, P = 0"

"Q= 300-50(0)"

"Q= 300-0= 300"